# The Early Bird Catches the Leak: Unveiling Timing Side Channels in LLM Serving Systems

### v4

@@ III-A Overview
Creating effective prompts is a challenging task that requires substantial effort, particularly in scenarios like in-context learning where extensive data is needed to optimize LLM performance. Furthermore, prompts can include personal or sensitive information, making them valuable assets that must be safeguarded. For instance, Samsung Electronics has prohibited employees from using generative AI tools like ChatGPT to prevent accidental disclosure of confidential data to OpenAI [46].
@@ III-B Prompt Stealing Attacks (PSA)
Background. KV caching is a widely adopted optimization in LLM serving systems, retaining the key and value embeddings from earlier inference steps to circumvent redundant computations during autoregressive text generation. Recent innovations, notably PagedAttention [14], improve on this concept by allowing the reuse of cached embeddings when prompts share common text segments, such as system messages, templates, or documents frequently included across multiple prompts. Representative implementations include automatic prefix sharing in vLLM [48], which detects shared prefix segments at runtime for KV cache reuse, and RadixAttention in SGLang [20], which efficiently manages shared KV caches for prompts containing common prefixes.
The side channel. Sharing KV caches can introduce a timing side channel. During the prefill phase of LLM inference, If a request’s prefix matches one already stored in the KV cache, it will be processed more quickly. Because most LLMs stream their outputs (i.e., token by token), it is possible to measure the fine-grained Time to First Token (TTFT) and detect timing discrepancies associated with cache hits versus misses. Assuming a stable network latency, we estimate the timing gap under a typical LLM deployment [49] as follows. Consider a model with 7 billion parameters (e.g., Llama-7B) running on an A100 GPU with 312 TFLOPS of computational power and a memory bandwidth of 1.5 TB/s. In the best case, with full GPU utilization, a cache miss for a single token during the prefill phase may take:
@@
These timing differences underpin the prompt stealing attack, which leverages KV cache sharing. Specifically, an attacker sends a request to the LLM and observes TTFT to detect whether the victim’s prefixes match. Since KV cache sharing occurs only for prompts with the same prefix, we devised an incremental search algorithm to recover prompts on a token-by-token basis. We present this algorithm and evaluate its real-world efficiency in Section IV-A.
@@ III-C Peeping Neighbor Attacks (PNA)
Background. Semantic caching (e.g., GPTCache [25, 28]) stores prior requests and their corresponding responses. Upon receiving a new request, it measures semantic similarity with cached requests. If the similarity surpasses a certain threshold, the system returns the cached response; otherwise, it queries the LLM again. Semantic caching can significantly reduce costs and enhance performance, and is integrated into major LLM frameworks like LangChain [27] and LlamaIndex [51].
The side channel. Unlike KV caching, which reuses data only for identical prompt prefixes, semantic caching allows reuse based on semantic similarity beyond a predefined threshold. However, sharing a semantic cache among multiple users can inadvertently reveal their requests. Cache hits provide responses in mere milliseconds, whereas cache misses can take several seconds—creating a clear timing difference that can be exploited by an attacker. This discrepancy enables the attacker to infer the semantics of concurrent requests issued by nearby users, a scenario we refer to as the peeping neighbor attack. Despite this, when the attacker tries to match a victim’s request semantically, the attacker’s own requests may also be cached, introducing noise into subsequent attempts. To address this challenge, we propose an efficient search algorithm that both minimizes the caching effects of the attacker’s own requests and improves the detection rate for the victim’s request. We describe this algorithm and illustrate how it can recover private information from neighboring users’ prompts in Section IV-B.
@@ IV-A Analysis on PSA
Attack setup. With increasing concerns about client data leakage in public LLM services, enterprise LLMs have become increasingly popular [52, 53]. In our study, we focus on a use case where the LLM API service is constructed from open-source projects within a local network environment. As described in Section II-D, we consider a scenario in which a victim develops a popular LLM application (e.g., a chatbot) using a proprietary system prompt via the LLM service. The attacker interacts with this LLM application over the local network, measures the TTFT, and attempts to uncover the system prompt based on timing discrepancies. Specifically, the LLM service uses the SGLang backend API server [26], which supports KV cache sharing for common prefixes. Notably, LLM API servers including OpenAI allow users to define various roles, such as “system” and “user”, within their requests [54], or send requests directly. Based on this capability, we categorize requests into two modes: synthesized and direct. The victim’s LLM chatbot, built on the FastChat framework [55], explicitly supports both, as shown in Figure 1. In direct mode, the user sends requests directly to the SGLang backend, whereas in synthesized mode, the full prompt is created by concatenating messages from each role according to predefined templates.
@@
Figure 3: Latency distribution of one token hit and miss. We used 2 representative models of different sizes: Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct-GPTQ-INT4. We used the Llama-3.1-70B-Instruct-GPTQ-INT4 model for subsequent evaluations.
@@
Characterizing the leakage. We began by examining the timing difference between a cache hit and a miss for a single token, comparing multiple model sizes. Specifically, we tested SGLang v0.3.0 [26], which organizes KV cache blocks in a radix tree to facilitate efficient prefix matching and reuse. Two models were evaluated: Llama-3.1-8B-Instruct [57] and Llama-3.1-70B-Instruct-GPTQ-INT4 [50]. System prompts of varying lengths were derived from the gabrielchua/system-prompt-leakage [58] dataset on Hugging Face. We measured TTFTs using 15-token prompts where only the last token differed, resulting in a shared-prefix length differing by exactly one—signifying cache hits versus misses—across 4,000 runs. Figure 3 illustrates the resulting time distributions, revealing a pronounced distinction between hit and miss scenarios for both models.
@@
Based on these observations, a straightforward classifier can be built to categorize the last token in the prompt as a hit if its latency is below a predefined threshold. However, as prompts grow longer, the latency also tends to rise due to increased computational demands, necessitating a distinct length-dependent threshold for each prompt length. To overcome this limitation, we measured the relative latency difference between hit and miss cases for last tokens across prompts of varying lengths (1–200 tokens). Our findings reveal that this difference is stable, independent of prompt length, and significantly larger than the variance within miss cases. This stability allows us to use a classifier with a single threshold instead of separate classifiers for different lengths, ensuring reliable hit detection.
@@
In real-world evaluations of our classifier, we noted that TTFT varies due to factors such as GPU system noise and power fluctuations, weakening the effectiveness of a fixed classification threshold, as shown in Figure 3. To mitigate this, we first construct a miss prompt by appending a rare token to a known prompt (i.e., requests without predicted tokens). Then, within a brief time window, we simultaneously collect TTFT data for both the miss and target prompts within a brief time window, using their difference for threshold-based classification.
@@
To illustrate the classifier’s effectiveness, consider an example using the Llama-3.1-70B-Instruct-GPTQ-INT4 [50] model under our evaluation settings. We build a classifier to determine whether the last token of prompts hits the KV cache. We test the classifier using 4,000 hits and 4,000 misses drawn from randomly selected system prompts, achieving a TPR of 0.88 and a FPR of 0.10. To further mitigate noise, we employ multi-sampling by collecting $n$ TTFT samples per token. The token is classified as a hit only if the number of hit detections exceeds a threshold $k$. With $n=10$ and $k=5$, the final classifier attains a TPR of 0.99 and an FPR of 0.003.
@@
End-to-end attacks. We built a local chatbot using the FastChat framework as the victim application. Its backend API server is configured with SGLang v0.3.0, which supports KV cache sharing for common prefixes. In this evaluation, we used the gabrielchua/system-prompt-leakage [58] dataset from Hugging Face, containing over 350,000 synthetic system prompts. We applied Llama-3.1-8B-Instruct [57] to compute embeddings for each prompt and visualized the results using Uniform Manifold Approximation and Projection (UMAP), which reduces dimensions of datasets and enables effective visualization of data distribution [59]. UMAP visualization confirmed substantial semantic diversity across the dataset, reflecting the varied topics and styles found in real-world prompt engineering. As shown in Figure 4, this semantic heterogeneity increases as the dataset expands, significantly complicating predictive modeling [60]. Consistent with the threat model detailed in Section II-D, we assume the attacker has a public dataset of system prompts that mirrors those of the victim. To more accurately reflect real-world conditions—-where production systems typically implement a small number of carefully crafted system prompts—-we randomly select only 200 prompts as the victim’s configuration while reserving the remainder as the attacker’s training corpus. This experimental design deliberately simulates the “needle in a haystack” challenge faced in realistic attack scenarios, where adversaries must identify specific target prompts within a substantially larger heterogeneous corpus.
@@
In constructing the next-token predictor, we adopt Llama-3.1-8B-Instruct [57] as base model, chosen as it shares an identical tokenizer with the victim LLM. We fine-tune this model using the attacker’s training dataset. Given the partially recovered prompt and the chosen temperature, the predictor generates a probability distribution for the next token. Temperature scaling introduces variability into the prediction. Our predictor is not heavily optimized; it was fine-tuned on a single NVIDIA A100 PCIE GPU for only a few hours with limited training resources. With larger models, more epochs, and bigger batch sizes, the predictor’s performance could be further improved.
To obtain TTFT values in the end-to-end scenario, the attacker first clears both the victim’s and the attacker’s requests from the cache. Next, the attacker measures the TTFT for their own request after the victim’s system prompt has been cached. Below, we describe how the TTFT is gathered and how caches are flushed.
$\bullet$ Timing measurement. As shown in Figure 5, the attacker begins by issuing a synthesized request containing the targeted system prompt. Once the end-of-sequence token is received in the POST response, a short delay is introduced, ensuring the system prompt resides in the KV cache. The attacker then sends a direct request via a POST call using the anticipated prompt, configured to generate only one token of output. The TTFT is computed as the interval between sending the request and detecting the first non-blank token in the streamed response.
@@
ttft = end_time - start_time
return ttft
@@
# Trigger the system prompt first
response = client.chat.completions
.create(...)
@@
$\bullet$ Flushing the caches through eviction. We observed SGLang provides a flush_cacheflush\_cache API [61] that efficiently clears the cache. However, for our end-to-end attack scenario, we chose not to use this API, as it is unlikely to be accessible to attackers in real-world environments. Instead, we employed a more robust method of evicting the KV cache by issuing batches of irrelevant requests. Under default SGLang settings, sending 15 such requests (each containing about 200 tokens) was sufficient to trigger eviction in about 5 seconds. This approach proved successful in 100% of our 10,000 tests.
@@
We evaluated both the token recovery accuracy and the average number of queries required per token. The results indicate a success rate of 89.0%, an FPR of 0.04, and an average of 5.57 guesses with 111.46 attack queries needed per recovered token when $n=10$. The corresponding cost per token recovery is approximately \$0.16 using OpenAI o1, \$0.012 with OpenAI o3-mini, and \$0.001 with Deepseek-Chat (as of March 5, 2025). Out of 200 victim prompts, we successfully recovered an average of 11.58 tokens for the top 100 prompts, and 17.9 tokens on average for the top 50 prompts. The maximum number of tokens recovered for a single prompt was 81, achieved with just 513 total guesses. Table II presents several examples of target system prompts alongside the prompts recovered via PSA. Notably, we limited the token prediction attempts to a maximum of 80 per position, ceasing further attempts if the correct token was not identified within these trials. Consequently, the primary constraint in reconstructing entire system prompts stems from the precision of the predictor rather than the accuracy of the classifier—an aspect that falls outside the scope of this study. In real-world attacks, the attacker can recover additional tokens by deploying a more advanced next-token predictor and increasing the maximum number of attack queries. A demo for the end-to-end attack is presented in our website [62].
TABLE II: Examples of recovered system prompts, including the number of attack queries and recovered tokens. We only listed travel planning-related prompts to demonstrate the PSA’s ability to recover diverse expressions.
No. Recovered system prompts #queries, #recovered tokens
1 In your role as a dedicated travel itinerary assistant, you will craft well-organized travel plans based on user preferences and interests. Gather information from the user’s inputs such as destinations, travel dates, and preferred activities to formulate a comprehensive itinerary. The output should include: dates, activities planned for each day, estimated costs, and important local information such as culture or tips. Emphasize clear, organized, $\cdots$ 10,260/81
2 You are programmed to function as a travel itinerary planner focusing exclusively on creating unique travel experiences. Provide tailored itineraries for destinations worldwide. $\cdots$ 3,620/26
3 Imagine you are a travel itinerary planner specializing in creating unique and personalized travel experiences. Your role is to craft itineraries that cater to the diverse interests and needs of travelers. $\cdots$ 1,360/32
@@ IV-B Analysis on PNA
Attack setup. In the PNA attack, we note that not all user queries contain sensitive data. An attacker is unlikely to pursue generic requests like “What’s the weather today?”. Instead, they are expected to focus on specific requests more likely to reveal private information. For example, a request such as “Draft a travel plan for my husband and me to Rome in September for a 3-day stay at the Hilton hotel” could expose personal and location details. By identifying such high-risk queries, the attacker can exploit timing side channels to recover private information from other users’ requests. For this purpose, the attacker could compile a list of privacy-related prompts from online sources (Table III). The attacker has obtained a list of private attributes, and its goal is to discover connections between entities and these private attributes, e.g., whether a specific user is linked to a particular medical condition.
Figure 7 illustrates the PNA steps. When semantic caching is used, the LLM stores victim requests and serves cached responses for similar queries. To exploit this channel, the attacker creates requests containing private attributes and monitors TTFT. A noticeable reduction in TTFT indicates that a cached victim’s request has been matched semantically with the attacker’s probe, thereby revealing private user data with high accuracy.
@@
TABLE III: Examples of user prompts that contain private attributes.
@@
In our experimental study, we chose LangChain as the LLM serving framework, given its widespread adoption and extensive application ecosystem [63]. The local chatbot system was built using LangChain and integrated with GPTCache [64] for semantic caching. We used gpt-3.5-turbo API as the backend LLM, and MedQuAD [65] as the evaluation dataset.
@@
Step 1. We randomly sampled 10 names from the Python package names-dataset [66] ($Names$). To obtain 10 random medical conditions ($Medconds$), we selected 10 semantically unrelated Q&A pairs from the MedQuAD dataset and used GPT-3.5-turbo to extract medical conditions from the questions. We then randomly chose 1 name ($Name_{0}$) and 1 medical condition ($Medcond_{0}$) as the private attributes to be recovered; the remaining pairs serve as the negative group (described below).
Step 2. Since real-world users may phrase the same content differently, we generated multiple sentences that share the same semantics as $T_{0}$. Specifically, we asked GPT-3.5-turbo to paraphrase $T_{0}$ into $n$ variations $\{T_{1},\dots,T_{n}\}$, each filled with $Name_{0}$ and $Medcond_{0}$.
Following these steps, we created two sample sets: (1) Positive Group: Sentences that are semantically similar to $T_{0}$, all containing $Name_{0}$ and $Medcond_{0}$. Formally, $\{(T_{i},Name_{0},Medcond_{0})\mid i=1,\dots,n\}$. (2) Negative Group: The original template $T_{0}$ populated with other names or medical conditions. Formally, $\{(T_{0},Name_{i},\\ Medcond_{j})\mid i\neq 0\text{ or }j\neq 0\}$.
@@
End-to-end attacks. We consider a typical open-world scenario in which a victim user requests healthcare assistance from an LLM. For instance, the user might submit a query with semantics similar to the template “compose a meeting agenda…” shown in Table III, but with various names and medical conditions. The user may also send queries unrelated to the targeted request. To simulate this, we model the user’s queries as follows:
$\bullet$ Type-1 (true samples): Queries with the specific name (e.g., “Alice”) and the specific medical condition (e.g., “heart disease”).
@@
$\bullet$ Type-4 (false samples): Queries unrelated to the target scenario.
We assume that the attacker focuses on uncovering the private attribute associations found in Type-1 queries. The victim can freely choose different paraphrases while preserving the same underlying semantics. To simulate this, we configure the victim to send five random requests per round: one Type-1 query (true sample) and four false samples (one Type-2, one Type-3, and two Type-4 requests). To measure the effectiveness of the attack, we use the TPR to assess how successfully the attacker retrieves private attributes from Type-1 requests. We also measure separate FPRs to capture how often the attacker incorrectly categorizes each of the three false sample types as positive.
To perform effective end-to-end attacks on the semantic cache, we must eliminate noise introduced by the attacker’s own requests, which also remain in the cache. To address this challenge, we developed a method that fully clears the semantic cache after each attack round. Specifically, our experiments show that under GPTCache’s default configurations, sending 1,000 semantically unrelated requests is sufficient to remove any leftover cache entries. In this scenario, we assume an attacker aims to determine whether a particular user (e.g., “Alice”) is associated with a specific medical condition (e.g., “heart disease”). Since the victim may use different phrases, the attacker issues multiple requests to enhance coverage and boost the TPR. However, increasing the total number of requests also raises the risk of false positives (i.e., a higher FPR), especially if new requests strongly resemble earlier ones. To mitigate this issue, the attacker prioritizes representative requests in the request space, thus increasing the overall number of queries while minimizing interference among them.

### v5

@@ III-A Overview
Creating effective prompts is a challenging task that requires substantial effort, particularly in scenarios like in-context learning where extensive data is needed to optimize LLM performance. Furthermore, prompts can include personal or sensitive information, making them valuable assets that must be safeguarded. For instance, Samsung Electronics has prohibited employees from using Chatgpt to prevent accidental disclosure of confidential data to OpenAI [25].
@@ III-B Prompt Stealing Attacks (PSA)
Background. KV caching is a widely adopted optimization in LLM which avoids redundant computation during autoregressive generation by storing key and value embeddings. Representative implementations include PagedAttention in vLLM [9], which detects shared prefix segments at runtime for KV cache reuse, and RadixAttention in SGLang [10], which efficiently manages shared KV caches in prefix tree.
The side channel. Sharing KV caches can introduce a timing side channel. In the prefill phase, if a request’s prefix matches one already stored in the KV cache, it will be processed more quickly. Since most LLMs stream their outputs (i.e., token by token), it is possible to measure the fine-grained TTFT and detect timing discrepancies associated with cache hits versus misses. Consider a model with 7 billion parameters (e.g., Llama-7B) running on an A100 GPU with 312 TFLOPS of computational power and a memory bandwidth of 1.5 TB/s. In the best case, with full GPU utilization, a cache miss for a single token during the prefill phase may take:
@@
These timing differences underpin the prompt stealing attack, which leverages KV cache sharing with the same prefix. After sending a request to the LLM, an attacker can observe TTFT to infer prefix matches. Building on this observation, we devised an incremental search algorithm to recover prompts on a token-by-token basis in Section IV-A.
@@ III-C Peeping Neighbor Attacks (PNA)
Background. Semantic caching (e.g., GPTCache [13]) stores prior requests and their corresponding responses. Upon receiving a new request, it measures semantic similarity with cached requests. If the similarity surpasses a certain threshold, the system returns the cached response; otherwise, it queries the LLM again. This technique reduces costs, improves performance, and is widely adopted in frameworks like LangChain [14] and LlamaIndex [27].
The side channel. Unlike KV caching, which reuses data only for identical prompt prefixes, semantic caching allows reuse based on semantic similarity beyond a predefined threshold. However, sharing a semantic cache among multiple users can inadvertently reveal their requests. Cache hits provide responses in mere milliseconds, whereas cache misses can take several seconds—creating exploitable timing differences.This discrepancy enables the attacker to infer the semantics of concurrent requests issued by nearby users, a scenario we refer to as the peeping neighbor attack. Despite this, when the attacker tries to match a victim’s request semantically, the attacker’s own requests may also be cached, introducing noise into subsequent attempts. To address this challenge, we propose an efficient search algorithm that both minimizes the caching effects of the attacker’s own requests and improves the detection rate for the victim’s request, detailed in Section IV-B.
@@ IV-A Analysis on PSA
Attack setup. With increasing concerns about client data leakage in public LLM services, local enterprise LLMs have become increasingly popular [28]. In our study, we focus on a use case where the LLM API service is constructed from open-source projects within a local network environment. As described in Section II-C, we consider a scenario in which a victim develops a popular LLM application (e.g., a chatbot) using a proprietary system prompt via the LLM service. The attacker interacts with this LLM application over the local network, measures the TTFT, and attempts to uncover the system prompt based on timing discrepancies. Specifically, the LLM service uses the SGLang backend API server [10], which supports KV cache sharing for common prefixes. Notably, LLM API servers including OpenAI allow users to define various roles, such as “system” and “user”, within their requests [29], or send requests directly. Based on this capability, we categorize requests into two modes: synthesized and direct. The victim’s LLM chatbot, built on the FastChat framework [30], explicitly supports both, as shown in Figure 1. In direct mode, the user sends requests directly to the SGLang backend, whereas in synthesized mode, the full prompt is created by concatenating messages from each role according to predefined templates.
@@
Figure 3: Latency distribution of one token hit and miss.
@@
Characterizing the leakage. We began by examining the timing difference between a cache hit and a miss for a single token across different model sizes. Specifically, we tested SGLang v0.3.0, which uses a radix tree to manage KV cache blocks for efficient prefix matching. Two models were evaluated: Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct-GPTQ-INT4. System prompts were derived from the gabrielchua/system-prompt-leakage [32] dataset. Using 15-token prompts differing only in the last token—ensuring a shared-prefix difference of exactly one—we collected TTFT measurements over 4,000 runs. Figure 3 illustrates the resulting time distributions, revealing a pronounced distinction between hit and miss scenarios for both models.
@@
Based on these observations, a simple classifier can be built to categorize the last token in the prompt as a hit if its latency is below a predefined threshold. However, as prompts grow longer, the latency also tends to rise due to increased computational demands, necessitating different thresholds for each prompt length. To address this, instead of the absolute latency, we measured the relative latency difference between hit and miss cases for last tokens across prompts of varying lengths (1–200 tokens). Our findings reveal that even though the absolute latency varies, this difference remains stable across prompt length, and significantly larger than the variance within miss cases, enabling a single unified threshold for reliable hit detection.
@@
In real-world evaluations of our classifier, we noted that TTFT varies due to factors such as GPU system noise and power fluctuations, weakening the effectiveness of a fixed classification threshold, as shown in Figure 3. To mitigate this, we first construct a “miss prompt” by appending a rare token to the partially recovered prompt (i.e., requests without predicted tokens) and designate the prompt being evaluated as hit or miss as the “candidate prompt”. Then, within a brief time window, we simultaneously collect TTFT data for both the miss and candidate prompts. By using the difference between these measurements rather than absolute values, we establish a threshold-based classification that remains robust against temporal system variations.
@@
To illustrate the classifier’s effectiveness, consider an example using the Llama-3.1-70B-Instruct-GPTQ-INT4 model under our evaluation settings. We built a classifier to determine whether the last token of prompts hit the KV cache. We tested the classifier using 4,000 hits and 4,000 misses drawn from randomly selected system prompts, achieving a TPR of 0.88 and an FPR of 0.10. To further mitigate noise, we employ multi-sampling by collecting $n$ TTFT samples per token. A token is classified as a hit only if the number of detections exceeds a threshold $k$. This leverages binomial statistics to amplify the TPR while suppressing the FPR. With $n=10$ and $k=5$, the final classifier attains a TPR of 0.99 and an FPR of 0.003.
@@
End-to-end attacks. We built a local chatbot using the FastChat as the victim application. Its backend API server is configured with SGLang v0.3.0, and the model deployed by victim is Llama-3.1-70B-Instruct-GPTQ-INT4. For evaluation, we used the gabrielchua/system-prompt-leakage dataset, containing over 350,000 synthetic system prompts. We applied Llama-3.1-8B-Instruct to compute embeddings for each prompt and visualized the results using Uniform Manifold Approximation and Projection (UMAP), which reduces the dimensions of datasets and enables effective visualization of data distribution [33]. UMAP visualization confirmed substantial semantic diversity across this dataset, reflecting the varied topics and styles found in real-world prompt engineering. As shown in Figure 4, this semantic heterogeneity increases as the dataset expands, significantly complicating predictive modeling. Consistent with the threat model detailed in Section II-C, we assume the attacker has a public dataset of system prompts that mirrors those of the victim. To reflect real-world conditions—where production systems typically implement a small number of carefully crafted system prompts—we randomly select 200 prompts as the victim’s configuration while reserving the remainder as the attacker’s training corpus. This experimental design deliberately simulates the “needle in a haystack” challenge faced in realistic attack scenarios, where adversaries must identify specific target prompts within a substantially larger heterogeneous corpus.
@@
In constructing the next-token predictor, we adopt Llama-3.1-8B-Instruct as the base model, chosen as it shares an identical tokenizer with the victim LLM. We fine-tune this model using the attacker’s training dataset. Given the partially recovered prompt and the chosen temperature, the predictor generates a probability distribution for the next token, while temperature scaling introduces variability into the prediction. Our predictor is not heavily optimized; it was fine-tuned on a single NVIDIA A100 PCIE GPU for only a few hours with limited training resources. With larger models, more epochs, and bigger batch sizes, the predictor’s performance could be further improved.
To obtain TTFT values, the attacker first clears both the victim’s and the attacker’s requests from the cache. Next, the attacker measures the TTFT for their own request after the victim’s system prompt has been cached. Below, we describe how the TTFT is gathered and how caches are flushed.
$\bullet$ Timing measurement. As shown in Figure 5, the attacker begins by issuing a synthesized request containing the targeted system prompt. Once the end-of-sequence token is received in the POST response, a short delay is introduced, ensuring the system prompt resides in the KV cache. The attacker then sends a direct request via a POST call using the predicted prompt, configured to generate only 1 output token. The TTFT is computed as the interval between sending the request and detecting the first non-blank token in the streamed response.
@@
return end_time - start_time
@@
response = requests.post(..)
@@
$\bullet$ Flushing the caches through eviction. We observed SGLang provides a flush_cacheflush\_cache API [34] that efficiently clears the cache. However, for our end-to-end attack scenario, we chose not to use this API, as it is unlikely to be accessible to attackers in real-world environments. Instead, we employed a more robust method of evicting the KV cache by issuing batches of irrelevant requests. These irrelevant requests consist of diverse texts with rare prefixes to rapidly fill the GPU memory and trigger cache eviction. Under default SGLang settings, sending 15 such requests (each containing about 200 tokens) was sufficient to trigger eviction in about 5 seconds. This approach proved successful in 100% of our 10,000 tests.
@@
We evaluated both the token recovery accuracy and the average number of queries required per token. The results indicate a success rate of 89.0%, an FPR of 0.04, and an average of 5.64 guesses with 112.72 attack queries needed per recovered token when $n=10$. The corresponding cost per token recovery is approximately \$0.16 using OpenAI o1, \$0.012 with OpenAI o3-mini, and \$0.001 with Deepseek-Chat (as of March 5, 2025). Out of 200 victim prompts, we successfully recovered an average of 11.32 tokens for the top 100 prompts, and 17.84 tokens on average for the top 50 prompts. The maximum number of tokens recovered for a single prompt was 81, achieved with just 519 total guesses. Notably, we limited the token prediction attempts to a maximum of 80 per position, ceasing further attempts if the correct token was not identified within these trials. Consequently, the primary constraint in reconstructing entire system prompts stems from the precision of the predictor rather than the accuracy of the classifier—an aspect that falls outside the scope of this study. In real-world attacks, the attacker can recover additional tokens by deploying a more advanced next-token predictor and increasing the maximum number of attack queries. A demo for the attack is presented on our website [35].
Discussions. Currently, we assume that the attacker has access to a public dataset of system prompts that can be used to fine-tune the next-token predictor. When a generic (i.e., untuned) predictor is employed, it tends to generate more false predictions, which in turn requires additional verification queries to distinguish genuine cache hits from prediction errors (i.e., to reduce false positives). This added overhead increases the effort and resources needed for an attacker to recover tokens. However, these challenges raise the cost of the attack rather than eliminating the underlying vulnerability. The core timing side channels we identified in KV-cache mechanisms remain exploitable regardless of predictor quality. With sufficient resources and persistence, attackers can still leverage these timing differences to recover tokens, albeit at a higher cost. To evaluate how predictor quality affects attack efficiency, we conducted the attack using the untuned Llama-3.1-8B-Instruct model as the generic predictor. Our results show a recovery rate of 69%, with an average of 163.48 guesses and 5,231.30 attack queries per token (here we set $n=16$ and $k=8$ to reduce false positives, and limit the token prediction attempts to a maximum of 1,000 per token), and an FPR of 1.5%. Further analysis reveals that, compared with fine-tuned predictors, generic predictors struggle to correctly predict certain rare tokens, leading to a lower recovery rate. Moreover, because generic predictors are less likely to quickly guess the target token, the average number of guesses rises sharply (from 5.64 to 163.48). This further demonstrates that our side-channel-based classifier is robust to false positives, as the attack maintains a low overall false positive rate (1.5%) even under significantly increased guessing attempts.
@@ IV-B Analysis on PNA
Attack setup. In the PNA attack, we observe that not all user queries contain sensitive data. An attacker typically ignores generic requests like “What’s the weather today?” and focuses on those specific revealing private information. For example, a prompt like “Draft a travel plan for me to Rome in September for a 3-day stay at the Hilton hotel” could expose personal and location details. We assume both the victim and the attacker are normal users of a chatbot service that uses a semantic cache in its backend. By identifying high-risk queries, the attacker can exploit timing side channels to recover private information from other users’ requests. For this purpose, the attacker could compile a list of privacy-related prompts from online sources (Table II). The attacker has obtained a list of private attributes, and aims to discover connections among these private attributes, e.g., whether a specific user is linked to a particular medical condition.
Figure 7 illustrates the PNA steps. When semantic caching is used, the LLM stores victim requests and serves cached responses for similar queries. To exploit this channel, the attacker creates requests containing private attributes and monitors TTFT. Since cache hits retrieve responses directly from local storage and bypass LLM inference, they produce a sharp TTFT reduction (from 5 seconds to under 1 second), which far exceeds the timing variations caused by network jitter. Consequently, by monitoring such timing differences, an attacker can infer semantic matches with a victim’s cached request and potentially expose private data.
@@
TABLE II: Examples of user prompts that contain private attributes.
@@
In our experimental study, we chose LangChain as the LLM serving framework, given its widespread adoption and extensive application ecosystem. The local chatbot system was built using LangChain and integrated with GPTCache for semantic caching. We used gpt-3.5-turbo API as the backend LLM, and MedQuAD [36] as the evaluation dataset. For evaluating semantic similarity, we employed $OnnxModelEvaluation$ method in GPTCache, which leverages the albert-duplicate-onnx model.
@@
Step 1. We randomly sampled 10 names from the Python package names-dataset ($Names$) and extracted 10 medical conditions ($Medconds$) from semantically unrelated Q&A pairs in MedQuAD dataset using GPT-3.5-turbo. We then randomly chose 1 name ($Name_{0}$) and 1 medical condition ($Medcond_{0}$) as the private attributes to be recovered; the remaining pairs serve as the negative group (described below).
Step 2. Since real-world users may phrase the same content differently, we generated multiple sentences that share the same semantics as $T_{0}$. Specifically, we asked GPT-3.5-turbo to paraphrase $T_{0}$ into $n$ variations $\{T_{1},\dots,T_{n}\}$, each filled with $Name_{0}$ and $Medcond_{0}$. Following these steps, we created two sample sets: (1) Positive Group: Sentences that are semantically similar to $T_{0}$, all containing $Name_{0}$ and $Medcond_{0}$. Formally, $\{(T_{i},Name_{0},Medcond_{0})\mid i=1,\dots,n\}$. (2) Negative Group: The original template $T_{0}$ populated with other names or medical conditions. Formally, $\{(T_{0},Name_{i},\\ Medcond_{j})\mid i\neq 0\text{ or }j\neq 0\}$.
@@
End-to-end attacks. We consider a typical open-world scenario in which a victim user requests healthcare assistance from an LLM. For instance, the user might submit a query with semantics similar to the template “compose a meeting agenda…” shown in Table II, but with various names and medical conditions. The user may also send queries unrelated to the targeted request. To simulate this, we model the user’s queries as follows:
$\bullet$ Type-1 (true samples): Queries with the specific name (e.g., “Alice”) and specific medical condition (e.g., “heart disease”).
@@
$\bullet$ Type-4 (false samples): Unrelated queries.
We assume that the attacker targets the private attribute associations in Type-1 queries. The victim can freely choose different paraphrases while preserving the same underlying semantics. To simulate this, we configure the victim to send five random requests per round: one Type-1 query (true sample) and four false samples (one Type-2, one Type-3, and two Type-4 requests). To measure the effectiveness of the attack, we use the TPR to assess how successfully the attacker retrieves private attributes from Type-1 requests. We also measure separate FPRs to capture how often the attacker misclassifies each of the three false sample types as positive.
To perform effective end-to-end attacks on the semantic cache, we must eliminate noise from the attacker’s own cached requests. We address this by clearing the cache after each attack round: under GPTCache’s default settings, sending 1,000 semantically unrelated requests suffices to remove any leftover cache entries. These requests were pre-extracted from multi-domain articles. Their semantic irrelevance was validated using SBERT [37], considering pairs with a similarity score below 0.3 as semantically unrelated. In this scenario, we assume an attacker aims to determine whether a particular user (e.g., “Alice”) is associated with a specific medical condition (e.g., “heart disease”). Since the victim may use different phrases, the attacker issues multiple requests to enhance coverage and boost the TPR. However, increasing the total number of requests also raises the risk of false positives, especially if new requests strongly resemble earlier ones. To mitigate this issue, the attacker prioritizes representative requests in the request space, thus increasing the overall number of queries while minimizing interference among them.